# Hutchinson Trace Estimation for High-Dimensional and High-Order Physics-Informed Neural Networks

### 22 Dec 2023 v1

@@ Abstract
Physics-Informed Neural Networks (PINNs) have proven effective in solving partial differential equations (PDEs), especially when some data are available by blending seamlessly data and physics. However, extending PINNs to high-dimensional and even high-order PDEs encounters significant challenges due to the computational cost associated with automatic differentiation in the residual loss function calculation. Herein, we address the limitations of PINNs in handling high-dimensional and high-order PDEs by introducing the Hutchinson Trace Estimation (HTE) method. Starting with the second-order high-dimensional PDEs, which are ubiquitous in scientific computing, HTE is applied to transform the calculation of the entire Hessian matrix into a Hessian vector product (HVP). This approach not only alleviates the computational bottleneck via Taylor-mode automatic differentiation but also significantly reduces memory consumption from the Hessian matrix to an HVP’s scalar output. We further showcase HTE’s convergence to the original PINN loss and its unbiased behavior under specific conditions. Comparisons with the Stochastic Dimension Gradient Descent (SDGD) highlight the distinct advantages of HTE, particularly in scenarios with significant variability and variance among dimensions. We further extend the application of HTE to higher-order and higher-dimensional PDEs, specifically addressing the biharmonic equation. By employing tensor-vector products (TVP), HTE efficiently computes the colossal tensor associated with the fourth-order high-dimensional biharmonic equation, saving memory and enabling rapid computation. The effectiveness of HTE is illustrated through experimental setups, demonstrating comparable convergence rates with SDGD under memory and speed constraints. Additionally, HTE proves valuable in accelerating the Gradient-Enhanced PINN (gPINN) version as well as the Biharmonic equation. Overall, HTE opens up a new capability in scientific machine learning for tackling high-order and high-dimensional PDEs.
@@ 2 Related Work
High-Dimensional PDE Solver. Several works have considered high-dimensional PDE solvers. In [51], the significance of $L^{\infty}$ loss and adversarial training in addressing high-dimensional Hamilton-Jacobi-Bellman equations was demonstrated. The separable PINNs approach in [12] employs a structure that allows the residual points to be a tensor product of per-dimension points, thereby increasing the batch size. Nevertheless, when confronted with problems surpassing ten dimensions, memory usage becomes a serious bottleneck. DeepBSDE [16, 17] and its extensions [2, 10, 19, 24, 31, 6] are grounded in the classical BSDE interpretation for specific high-dimensional parabolic PDEs, leveraging deep learning models to approximate the unknowns in the BSDE formulation. The deep splitting method [1] unifies the splitting method with deep neural networks. Chen et al. [11] solved forward and inverse problems of Fokker-Planck equations, including Brownian noise and Levy noise in high dimensions using PINNs. FBSNN [44] established a connection between high-dimensional parabolic PDEs and forward-backward stochastic differential equations, employing deep neural networks to learn the unknown solution. The multilevel Picard methods [3, 4, 5, 26, 27] represent a nonlinear extension of Monte Carlo capable of solving parabolic PDEs under specific constraints. In [52, 53], tensor neural networks were proposed, adopting a separable structure for efficient numerical integration in solving high-dimensional Schr”odinger equations. More recently, SDGD [22] emerged as a method designed to sample dimensions in PDEs, aiming to scale up and accelerate high-dimensional PINNs. Randomized smoothing [18, 23] adopts the smoothed model with Gaussian noise so that Monte Carlo can simulate its inference and derivatives to combat the curse of dimensionality and avoid the costly automatic differentiation in high-order and high dimensions.
@@ 3.1 Preliminaries
Given the boundary points $\{\boldsymbol{x}_{b,i}\}_{i=1}^{n_{b}}\subset\Gamma$ and the residual points $\{\boldsymbol{x}_{r,i}\}_{i=1}^{n_{r}}\subset\Omega$, PINNs minimize the discrepancy in the residual and on the boundary:
@@ 3.2 HTE for High-Dimensional Second-Order Parabolic PDEs
where $A_{\theta}(\boldsymbol{x},t):=\sigma\sigma^{\mathrm{T}}(\boldsymbol{x},t)\left(\operatorname{Hess}_{\boldsymbol{x}}u\right)(\boldsymbol{x},t)\in\mathbb{R}^{d\times d}$ and $B_{\theta}(\boldsymbol{x},t):=\partial_{t}u(\boldsymbol{x},t)+f(\boldsymbol{x},t,u,\nabla_{\boldsymbol{x}}u)\in\mathbb{R}.$
@@
The proof is presented in Appendix A.1. ∎
@@
To conclude this subsection, we provide an implementation of HTE in a few lines of code using JAX [8]. Although HTE is mathematically formulated as an HVP, the implementation details significantly impact its speed and memory usage. For instance, a straightforward computation of the entire Hessian followed by a vector multiplication will not yield any efficiency gains. Below, we show a sample code for highly efficient HTE in JAX Taylor-mode autodiff, based on fully forward mode [7], which is super fast due to the absence of backward mode autodiff.
@@
3.3 Comparison with SDGD: Variance and Applicability
Recently, the related work of SDGD [22] has demonstrated robust capabilities in tackling high-dimensional PDEs and the curse-of-dimensionality. SDGD primarily relies on sampling dimensions to mitigate the computational and memory burdens associated with increasing dimensions. In contrast, HTE employs HVP to avoid the challenges posed by computing the full Hessian, whose cost grows quadratically with dimensionality. In this subsection, we will compare their variances and discuss their similarity and applicability.
@@ 3.3.2 Variance Comparison between HTE and SDGD
3.3.3 Applicability Comparison between HTE and SDGD
Next, we move to discuss the applicability of HTE compared to SDGD. HTE is more beneficial in higher dimensions and higher orders; otherwise, regular autograd methods may already be sufficient. For instance, consider the extreme case of one dimension. In this scenario, the Hessian is just a scalar. Consequently, the computational cost and memory consumption for both HVP and the full Hessian are exactly the same, hence providing no advantage for HTE. Another example will be high-dimensional but low-order; while the HTE approach can be applied to first-order PDEs, i.e., the norm square of the Jacobian can be unbiasedly estimated by Jacobian-vector product (JVP):
$\|\nabla_{\boldsymbol{x}}u(\boldsymbol{x})\|^{2}=\mathbb{E}_{v\sim p(v)}|v^{\mathrm{T}}\nabla_{\boldsymbol{x}}u(\boldsymbol{x})|^{2},$ (12)
experiments indicate that memory savings are insignificant. This term $\|\nabla_{\boldsymbol{x}}u(\boldsymbol{x})\|^{2}$ is ubiquitous in scientific computing, e.g., the first-order Hamilton-Jacobi equation, and the Schrödinger equation in the variation form. (see section 5.4 of SDGD [22] where SDGD scales up tensor neural networks for the Schrödinger equation).
More generally, from a mathematical perspective, for a $d$-dimensional $n$th-order PDE, the pure autograd computation results in a tensor of dimensions $d^{n}$. For example, the Jacobian is of dimension $d$, and the Hessian is of dimension $d^{2}$. In contrast, the output of HTE are $V$ scalars when using $V$ as the batch size; thus, the output dimension is $V$. Consequently, as the dimension and order increase, the advantages of HTE become more pronounced.

### 03 Mar 2024 v2

@@ Abstract
Physics-Informed Neural Networks (PINNs) have proven effective in solving partial differential equations (PDEs), especially when some data are available by seamlessly blending data and physics. However, extending PINNs to high-dimensional and even high-order PDEs encounters significant challenges due to the computational cost associated with automatic differentiation in the residual loss function calculation. Herein, we address the limitations of PINNs in handling high-dimensional and high-order PDEs by introducing the Hutchinson Trace Estimation (HTE) method. Starting with the second-order high-dimensional PDEs, which are ubiquitous in scientific computing, HTE is applied to transform the calculation of the entire Hessian matrix into a Hessian vector product (HVP). This approach not only alleviates the computational bottleneck via Taylor-mode automatic differentiation but also significantly reduces memory consumption from the Hessian matrix to an HVP’s scalar output. We further showcase HTE’s convergence to the original PINN loss and its unbiased behavior under specific conditions. Comparisons with the Stochastic Dimension Gradient Descent (SDGD) highlight the distinct advantages of HTE, particularly in scenarios with significant variability and variance among dimensions. We further extend the application of HTE to higher-order and higher-dimensional PDEs, specifically addressing the biharmonic equation. By employing tensor-vector products (TVP), HTE efficiently computes the colossal tensor associated with the fourth-order high-dimensional biharmonic equation, saving memory and enabling rapid computation. The effectiveness of HTE is illustrated through experimental setups, demonstrating comparable convergence rates with SDGD under memory and speed constraints. Additionally, HTE proves valuable in accelerating the Gradient-Enhanced PINN (gPINN) version as well as the Biharmonic equation. Overall, HTE opens up a new capability in scientific machine learning for tackling high-order and high-dimensional PDEs.
@@ 2 Related Work
High-Dimensional PDE Solver. Several works have considered high-dimensional PDE solvers. In [53], the significance of $L^{\infty}$ loss and adversarial training in addressing high-dimensional Hamilton-Jacobi-Bellman equations was demonstrated. The separable PINNs approach in [12] employs a structure that allows the residual points to be a tensor product of per-dimension points, thereby increasing the batch size. Nevertheless, when confronted with problems surpassing ten dimensions, memory usage becomes a serious bottleneck. DeepBSDE [16, 17] and its extensions [2, 10, 19, 25, 32, 6] are grounded in the classical BSDE interpretation for specific high-dimensional parabolic PDEs, leveraging deep learning models to approximate the unknowns in the BSDE formulation. The deep splitting method [1] unifies the splitting method with deep neural networks. Chen et al. [11] solved forward and inverse problems of Fokker-Planck equations, including Brownian noise and Levy noise in high dimensions using PINNs. FBSNN [45] established a connection between high-dimensional parabolic PDEs and forward-backward stochastic differential equations, employing deep neural networks to learn the unknown solution. The multilevel Picard methods [3, 4, 5, 27, 28] represent a nonlinear extension of Monte Carlo capable of solving parabolic PDEs under specific constraints. The deep Galerkin method (DGM) [49] trains neural networks to satisfy the high-dimensional PDE operators and other conditions, where derivatives are estimated via Monte Carlo. In [54, 55], tensor neural networks were proposed, adopting a separable structure for efficient numerical integration in solving high-dimensional Schr”odinger equations. More recently, SDGD [22] emerged as a method designed to sample dimensions in PDEs, aiming to scale up and accelerate high-dimensional PINNs. Randomized smoothing [18, 23] adopts the smoothed model with Gaussian noise so that Monte Carlo can simulate its inference and derivatives to combat the curse of dimensionality and avoid the costly automatic differentiation in high-order and high dimensions. HTE is also used in scale up and speed up score-based diffusion models [51, 24].
@@ 3.1 Preliminaries
Given the boundary points $\{\boldsymbol{x}_{b,i}\}_{i=1}^{n_{b}}\subset\Gamma$ and the residual points $\{\boldsymbol{x}_{r,i}\}_{i=1}^{n_{r}}\subset\Omega$, PINNs minimize the discrepancy in the residual and on the boundary for the PINN neural network model $u_{\theta}$ parameterized by $\theta$:
@@ 3.2.1 Methodology
where $A_{\theta}(\boldsymbol{x},t):=\sigma\sigma^{\mathrm{T}}(\boldsymbol{x},t)\left(\operatorname{Hess}_{\boldsymbol{x}}u_{\theta}\right)(\boldsymbol{x},t)\in\mathbb{R}^{d\times d}$, $B_{\theta}(\boldsymbol{x},t):=\partial_{t}u_{\theta}(\boldsymbol{x},t)+f(\boldsymbol{x},t,u_{\theta},\nabla_{\boldsymbol{x}}u_{\theta})\in\mathbb{R}$, and $u_{\theta}(\boldsymbol{x},t)$ is the PINN neural network model parameterized by $\theta$.
@@
The proof is presented in A.1. ∎
3.2.2 Discussion on the Bias and Unbiased HTE Versions
@@
Regarding the theoretical understanding of the biased version of HTE, this bias is proportional to the HTE-PINN residual variance. Concretely, its bias against the exact PINN loss $L_{\text{PINN}}(\theta)$ can be computed as follows:
$\quad\mathbb{E}_{\{\boldsymbol{v}_{i}\}_{i=1}^{V}}\left[L_{\text{HTE}}(\theta;\{\boldsymbol{v}_{i}\}_{i=1}^{V})\right]-L_{\text{PINN}}(\theta)$ (11)
$=\frac{1}{2}\mathbb{E}_{\{\boldsymbol{v}_{i}\}_{i=1}^{V}}\left[\left(\frac{1}{V}\sum_{i=1}^{V}\boldsymbol{v}_{i}^{\mathrm{T}}A_{\theta}(\boldsymbol{x},t)\boldsymbol{v}_{i}+B_{\theta}(\boldsymbol{x},t)\right)^{2}\right]-\frac{1}{2}\left[\operatorname{Tr}\left(A_{\theta}(\boldsymbol{x},t)\right)+B_{\theta}(\boldsymbol{x},t)\right]^{2}$
$=\frac{1}{2}\mathbb{E}_{\{\boldsymbol{v}_{i}\}_{i=1}^{V}}\left[\left(\frac{1}{V}\sum_{i=1}^{V}\boldsymbol{v}_{i}^{\mathrm{T}}A_{\theta}(\boldsymbol{x},t)\boldsymbol{v}_{i}+B_{\theta}(\boldsymbol{x},t)\right)^{2}\right]-\frac{1}{2}\left(\mathbb{E}_{\{\boldsymbol{v}_{i}\}_{i=1}^{V}}\left[\frac{1}{V}\sum_{i=1}^{V}\boldsymbol{v}_{i}^{\mathrm{T}}A_{\theta}(\boldsymbol{x},t)\boldsymbol{v}_{i}+B_{\theta}(\boldsymbol{x},t)\right]\right)^{2}$
$=\frac{1}{2}\mathbb{V}_{\{\boldsymbol{v}_{i}\}_{i=1}^{V}}\left[\frac{1}{V}\sum_{i=1}^{V}\boldsymbol{v}_{i}^{\mathrm{T}}A_{\theta}(\boldsymbol{x},t)\boldsymbol{v}_{i}+B_{\theta}(\boldsymbol{x},t)\right].$
which is the variance of the HTE-PINN residual defined as $\frac{1}{V}\sum_{i=1}^{V}\boldsymbol{v}_{i}^{\mathrm{T}}A_{\theta}(\boldsymbol{x},t)\boldsymbol{v}_{i}+B_{\theta}(\boldsymbol{x},t)$, which replaces the full PINN residual by HTE. Suppose learning rate decay is used (all our experiments adopt this common practice in machine learning). In that case, the scale of the loss for PINN optimization will become smaller, i.e., we are basically optimizing $\epsilon^{2}\cdot L_{\text{HTE}}(\theta;\{\boldsymbol{v}_{i}\}_{i=1}^{V})$ with a small and decaying $\epsilon>0$ in practice. Thus, the biased version of HTE’s bias becomes $\epsilon$ times the HTE-PINN residual variance, i.e. $\epsilon\times\frac{1}{2}\mathbb{V}_{\{\boldsymbol{v}_{i}\}_{i=1}^{V}}\left[\frac{1}{V}\sum_{i=1}^{V}\boldsymbol{v}_{i}^{\mathrm{T}}A_{\theta}(\boldsymbol{x},t)\boldsymbol{v}_{i}+B_{\theta}(\boldsymbol{x},t)\right]$, which decays gradually due to decaying $\epsilon$ to ensure a decreasing variance for better convergence.
In practice, the success of the biased version should be attributed to its relatively low bias. I.e., The biased version is the fastest due to sampling only once, and it can converge well thanks to the low bias in practice, as long as its batch size $V$ in HTE is not too small.
The corresponding experiments for the bias and unbiasedness are conducted in Section 4.1.2.
3.2.3 HTE Implementation Detail
We provide an implementation of HTE in a few lines of code using JAX [8]. Although HTE is mathematically formulated as an HVP, the implementation details significantly impact its speed and memory usage. For instance, a straightforward computation of the entire Hessian followed by a vector multiplication will not yield any efficiency gains, where the curse of dimensionality would still plague as the computation complexity scales with the PDE problem’s dimensionality. This is because we will need to compute all the Hessian trace elements using the standard basis and the number of HVP to compute scales quadratically with the dimension of the equation. The key reason that HTE+HVP works is that, instead of computing the exact hessian diagonal, we just compute a stochastic approximation of it, and the computation cost is amortized over the optimization process. With HTE, the number of HVPs to compute is constant regarding PDE dimensionality. Hence, the computation complexity scales much slower. Below, we show a sample code for highly efficient HTE in JAX Taylor-mode automatic differentiation (AD), based on fully forward mode [7], which is super fast due to the absence of backward mode autodiff. The algorithm of HTE aligns well with efficient Taylor mode AD to avoid the slow naive implementation; thus, it is employed to implement HTE. At the time of writing, JAX is the only mature AD framework that provides Taylor mode AD, so we chose to implement our method in JAX.
@@ 3.2.3 HTE Implementation Detail
Our approach and code implementation alleviate the cost of computing high-order derivatives by combining Taylor mode automatic differentiation (AD) and HTE, whose scaling is much better than the vanilla backward AD approaches widely adopted in the PINN literature.
In the AD framework, functions are implemented as the composition of primitives, which are simple functions with known analytical formulas for derivatives. We can write such function as $f=f_{n}\circ\dots\circ f_{1}$ where $f_{i}$ are the primitives, e.g., a neural network containing multiple linear layers and activations. Then by chain rule, the Jacobian of $f$ is simply the matrix product of the Jacobians of all the primitives $f_{i}$: $J_{f}=J_{f_{n}}\cdot\dots\cdot J_{f_{1}}$. Instead of computing the full Jacobians, in the AD framework, we usually compute a certain contraction of the derivative tensor, and it is computed by composing the analytical formula of the contraction considered for each primitive. The forward mode AD computes the Jacobi-Vector-Product (JVP): $J_{f}\boldsymbol{v}=J_{f_{n}}(J_{f_{n-1}}(\dots(J_{f_{1}}\boldsymbol{v})))$, whereas the backward mode AD computes the Vector-Jacobian-Product (VJP): $\boldsymbol{v}^{\mathrm{T}}J_{f}=(((\boldsymbol{v}^{\mathrm{T}}J_{f_{n}})J_{f_{n-1}})\dots J_{f_{1}})$. The vector $\boldsymbol{v}$ is called tangent in the case of JVP and cotangent in the case of VJP. From the perspective of differential geometry, the JVP of each primitive can be seen as the pushforward map from the input space to the next hidden space; hence, in AD, we also refer to the computation process of iteratively applying the JVP map of each primitive as “pushing the tangent forward”. Similarly, VJP can be seen as the pullback of the input cotangent $\boldsymbol{v}^{\mathrm{T}}$.
In the backward mode, one needs to first perform one forward pass to obtain the linearization points of all the primitive Jacobians, and this set of linearization points, also called the evaluation trace, needs to be stored in the memory. Furthermore, due to the forward pass, the number of sequential computations needed with the backward mode is $2n$ where $n$ is the number of primitives $\{f_{i}\}_{i=1}^{n}$, whereas for the forward mode, it is just $n$. This increases the computational complexity as well.
However, backward mode AD is the default choice for deep learning and PINN. This is because the Jacobian of a scalar loss function has the shape $1\times N$ where $N$ is the input dimension, which can be computed with one VJP, or $N$ JVP. Obviously, it would be much more efficient to do one VJP instead of $N$ JVP, even though for one contraction, VJP is more expensive than JVP in both speed and memory.
Therefore, the most widely adopted way in the PINN community to compute higher-order derivatives in the residual loss is to use repeated application of backward mode AD. But with each repeated application of backward mode AD, one doubles the amount of sequential computation and also the number of tensors to store in the evaluation trace for the backward pass. Furthermore, the new element to store in the evaluation trace has a dimension that grows exponentially to the derivative order. In this case, computing high-order derivatives with standard backward AD is indeed inefficient. Repeated application of forward mode AD is also inefficient, as the size of the input tangent $\boldsymbol{v}$ grows exponentially to the derivative order.
In this paper, we use Taylor mode AD, a generalization of forward mode AD that provides composition rules for a specific set of higher-order derivative tensor contractions related to the high-order Taylor expansion. In forward mode AD, only the JVP contraction is pushed forward, but in Taylor mode AD, we are pushing forward multiple contractions of derivatives tensor up to the highest orders. This avoids repeated application of AD when computing high-order derivatives. And just like forward mode AD, the huge derivative tensor was never materialized in Taylor mode AD, and only the contraction (a scalar) was computed. The memory complexity is, therefore, much lower than backward mode AD since no evaluation trace is required, and only the contracted quantities (a scalar) are stored in memory. Furthermore, since the number of random vectors $V$ in the HTE is fixed and unrelated to the dimension, the curse of dimensionality is also alleviated.
3.3 Comparison with SDGD
Recently, the related work of SDGD [22] has demonstrated robust capabilities in tackling high-dimensional PDEs and the curse-of-dimensionality. SDGD primarily relies on sampling dimensions to mitigate the computational and memory burdens associated with increasing dimensions. In contrast, HTE employs HVP to avoid the challenges posed by computing the full Hessian, whose cost grows quadratically with dimensionality. In this subsection, we will compare their variances and discuss their similarity.
@@ 3.3.1 SDGD as a Special Case of HTE
Overall, HTE and SDGD are similar in some aspects but differ in other viewpoints.
@@ 3.4 Applications to Biharmonic Equations
3.5 Applicability of HTE
Next, we move to discuss the applicability of HTE. HTE is more beneficial in higher dimensions and higher orders; otherwise, HTE’s acceleration effect will be less obvious, although it is still faster. More generally, from a mathematical perspective, for a $d$-dimensional $n$th-order PDE, the pure autograd computation results in a tensor of dimensions $d^{n}$. For example, the Jacobian is of dimension $d$, and the Hessian is of dimension $d^{2}$. In contrast, the output of HTE are $V$ scalars when using $V$ as the batch size; thus, the output dimension is $V$. Consequently, as the dimension and order increase, the advantages of HTE become more pronounced.
3.5.1 HTE’s Application to Other Neural PDE Solvers
Furthermore, our HTE approach covers various indispensable PDE differential operators. Hence, our HTE can be applied to other deep learning-based PDE solvers, such as the deep Ritz [56] and weak adversarial network (WAN) approaches [59] in addition to PINN [46].
Specifically, the deep Ritz method [56] considers variational problems with the differential operator $\|\nabla_{\boldsymbol{x}}u\|^{2}$, which can be efficiently estimated via $\|\nabla_{\boldsymbol{x}}u(\boldsymbol{x})\|^{2}=\mathbb{E}_{\boldsymbol{v}\sim p(\boldsymbol{v})}|\boldsymbol{v}^{\mathrm{T}}\nabla_{\boldsymbol{x}}u(\boldsymbol{x})|^{2},$ where $\mathbb{E}_{\boldsymbol{v}\sim p(\boldsymbol{v})}[\boldsymbol{v}\boldsymbol{v}^{\mathrm{T}}]=I$ as in our HTE method. The randomization of our HTE simplifies the original $O(d)$ ($d$ is the problem dimension) neural network gradients with respect to $\boldsymbol{x}\in\mathbb{R}^{d}$ into $O(1)$ Jacobian vector products (JVPs), thereby conserving memory consumption and accelerating computation. Here, the JVP employed is essentially a specific instance of the proposed Tensor-Vector Product (TVP) in our HTE method, where the tensor corresponds to the Jacobian.
Besides, Weak Adversarial Networks (WANs) [59] consider PDE weak formulations, which consider the second-order elliptic PDEs and second-order parabolic PDEs. These PDEs considered by WAN are a subset of the cases in our HTE paper with the following form: $\partial_{t}u(\boldsymbol{x},t)+{\operatorname{Tr}\left(\sigma\sigma^{\mathrm{T}}(\boldsymbol{x},t)\left(\operatorname{Hess}_{\boldsymbol{x}}u\right)(\boldsymbol{x},t)\right)}+f(\boldsymbol{x},t,u,\nabla_{\boldsymbol{x}}u)=0,\boldsymbol{x}\in\mathbb{R}^{d},t\in[0,T],$ Due to the same reason, our HTE can also be adopted to accelerate the convergence of WAN. In summary, our HTE serves as an acceleration technique for general high-dimensional and high-order neural PDE solvers, making it widely applicable in PINN, deep Ritz method, and WAN.
3.5.2 Discussion on HTE’s Applicable PDE Types
Furthermore, regarding the types of PDEs where HTE can be applied, our primary focus is on high-order and high-dimensional PDEs, where high-dimensional PDEs mainly consist of second-order elliptic and parabolic equations, which have wide-ranging applications [1, 16, 22, 23, 45], which covers the Fokker-Planck equation from statistical mechanics to describe the evolution of stochastic differential equations (SDEs), the Black-Scholes equation in mathematical finance, the Hamilton-Jacobi-Bellman equation in optimal control, and Schrödinger equation in quantum physics. Although it is possible to design irregular equations that render HTE unusable, such equations lack practical applications, and even the uniqueness of their solutions is not theoretically guaranteed. To the best of our knowledge, the high-dimensional PDEs HTE can solve can cover the majority of real-world PDEs with multiple applications. Additionally, for high-order equations, we consider high-order and high-dimensional biharmonic equations modeling elastic membranes, thin plates, and stream functions in fluid dynamics.
3.5.3 Discussion on Low-Dimensional and High-Order PDEs
Our HTE directly applies to high-order KdV and Kuramoto-Sivashinsky equations thanks to the efficient Taylor-mode automatic differentiation in HTE. More specifically, our HTE and the provided pseudocode can be directly applied to these high-order equations, with the distribution of the vector $\boldsymbol{v}$ being Rademacher distribution. Just to note that these equations are one-dimensional in $\boldsymbol{x}$, which is low-dimensional. In the case of one-dimensional but higher-order scenarios, the primary acceleration of HTE stems from Taylor-mode automatic differentiation, which, as previously mentioned, is significantly faster than the backward mode and forward mode conventionally employed in deep learning and PINN. The acceleration provided by our HTE becomes more pronounced as the dimensionality and order of the PDE problem increase. It often enables handling challenging cases where standard PINN would exceed GPU memory capacity or converge extremely slowly.